OpenStudy (blockcolder):

Easy brain teaser: Evaluate \[\sum_{k=0}^n \frac{k}{(k+1)!}\]

OpenStudy (unklerhaukus):

\[=0+\frac12+\frac26+\frac3{24}+\frac4{120}+\cdots\]\[=0+\frac12+\frac13+\frac18+\frac1{30}+\cdots\]

OpenStudy (unklerhaukus):

\[\approx0+0.5+0.333+0.125+0.033+\cdots\]\[=0.5+0.457+0.033\]\[=0.5+0.490\]\[=0.99\]\[\approx1\]

OpenStudy (blockcolder):

It's not until infinity. :P

OpenStudy (auctoratrox):

you add consecutive terms until you get + n/(n+1)! so it's 0 + ... + n/(n+1)! the number of terms you want to add depends on your preference.

OpenStudy (anonymous):

k=k+1-1 in num can help you.

OpenStudy (blockcolder):

I want the answer to not have a summation. (Of course, I know the answer.)

OpenStudy (anonymous):

sig(1/k!-1/k+1!)=1-0=1

OpenStudy (blockcolder):

The answer I wanted was \(1-\frac{1}{(n+1)!}\) but this will do.

OpenStudy (anonymous):

n is going to inf then i/n+1! going to 0

OpenStudy (anonymous):

ok i thought uper was inf

OpenStudy (anonymous):

yeap 1-1/n+1! is answer

OpenStudy (anonymous):

dont forget k=k+1-1